Please help me figure out how to graph and find the plots for each consultants equation. Here is what you need to know:
The coach of the Women's Olympic Gymnastic Team has a difficult decision to make. A strong contender for the team suffered a sprained knee on July 20, just about one month prior to the Olympic tryouts, and it on the disabled list. It is now just over two weeks later, August 6. The coach needs to decide whether or not to keep the gymnast on the disabled list.
7/20: Miss Tumble was administered a Cybex strength test upon arrival.Her quadriceps of the injured leg measured 55 foot-pounds in extension.
7/25: Daily regimen is showing progress. Cybex test measures 90 foot-pounds.
8/1: Some swelling earlier in the week.General reports of less pain. Cybex test measures 140 foot-pounds.
8/6: Less swelling. Range of motion has shown marked increase. Cybex test measures 185 foot-pounds.
a. Please graph the data above provided by the physical therapist.Let July 20 be represented as Day 1,which makes August 18 Day 30.
Two consultants were hired to advise the coach.
Consultant A said that the function F given by the equation f(x)=40+8x approximated the data well. So consultant A predicted that on August 18(Day 30), Miss Tumble would have 280 foot-pounds of strength and would be strong enough to compete.
Consultant B said the function g given by the equation g(x)=55+6x approximated the data well. So consultant B predicted that on August 18(Day 30), Miss Tumble would have only 235 foot-pounds of strength and would not be strong enough to compete.
b.Sketch each consultant's function on the graph you created in question a above. Use a different color pencil or pen for each graph.
CAN YOU PLEASE HELP ME FIGURE OUT HOW TO FIND THE COORDINATES FOR EACH CONSULTANT USING THEIR EQUATIONS AND HOW DO I GRAPH THEM? I am REALLY getting FRUSTRATED with algebra and I'm upset and confused. Can you PLEASE HELP ME???
Signed,
Confused and Frustrated Algebra Student

Answer by stanbon(48574)

7/20: Miss Tumble was administered a Cybex strength test upon arrival.Her quadriceps of the injured leg measured 55 foot-pounds in extension.
7/25: Daily regimen is showing progress. Cybex test measures 90 foot-pounds.
8/1: Some swelling earlier in the week.General reports of less pain. Cybex test measures 140 foot-pounds.
8/6: Less swelling. Range of motion has shown marked increase. Cybex test measures 185 foot-pounds.
---------
This data gives you four points (day,foot-pound)
The points are (1,55),(6,90),(13,140),(18,185)
Using the two middle points you can find the equation relating
day and foot-pounds:
slope= [140-90]/[13-6]=50/7
y=intercept: 140=(50/7)(13)+b; b=47.14
EQUATION: y=(50/7)x+47.14
Comment: The four points don't all lie on the same line. I picked
the two middle points. You might have picked two others.
-----------------
a. Please graph the data above provided by the physical therapist.Let July 20 be represented as Day 1,which makes August 18 Day 30.
Comment: I tried to show you the graph but this software will not display
the line.
----------------
Two consultants were hired to advise the coach.
Consultant A said that the function F given by the equation f(x)=40+8x approximated the data well. So consultant A predicted that on August 18(Day 30), Miss Tumble would have 280 foot-pounds of strength and would be strong enough to compete.
Consultant B said the function g given by the equation g(x)=55+6x approximated the data well. So consultant B predicted that on August 18(Day 30), Miss Tumble would have only 235 foot-pounds of strength and would not be strong enough to compete.
b.Sketch each consultant's function on the graph you created in question a above. Use a different color pencil or pen for each graph.
Comment: Again I tried to show you the two lines but to no avail.
Hope this helps.
Cheers,
Stan H.